choose some values of x and evaluate the expression for those values
1/(x-5)>10

Answer by CharlesG2(828)

choose some values of x and evaluate the expression for those values
1/(x-5)>10
if 1/(x-5) = 10 what would x be?
1 = 10 * (x-5) (multiplied both sides by x-5)
1 = 10x - 50
51 = 10x
5.1 = x
if 1/(x-5) = 20 what would x be?
1 = 20 * (x-5) (multiplied both sides by x-5)
1 = 20x - 100
101 = 20x
5.05 = x
x 1/(x-5) is >10 ?
0 -0.2 no
4 -1 no
4.9 -10 no
4.99 -100 no
5 #DIV/0! #DIV/0!
5.01 100 yes
5.02 50 yes
5.03 33.33333333 yes
5.04 25 yes
5.05 20 yes
5.09 11.11111111 yes
5.1 10 no
5.2 5 no
5.3 3.333333333 no
5.4 2.5 no
5.5 2 no
6 1 no
5 < x < 5.1